# A Review of Hydrogen-Enabled Resilience Enhancement for Multi-Energy Systems

### v2

@@ IV-B Definitions of HMES Resilience
Before describing the definitions of HMES resilience, we first provide two system performance curves related to MES and HMES under extreme events in Fig. 5, respectively. Although the typical curve under MES cannot perfectly capture the realistic system performance trajectories (e.g., local performance fluctuations) as illustrated in [68, 69, 70], it can reflect the system’s ability to prepare for, absorb, and recover from disturbances and has been widely used in existing works[50],[71, 72, 73, 74, 75],[115]. In addition, While it is not tailored to hydrogen-enabled MESs, the performance trajectory can be applied to hydrogen-enabled MESs after minor modifications. By comparing the differences between two curves, the physical and functional features of hydrogen can be observed. Firstly, we introduce the whole process of resilience enhancement under MES. In existing works, load loss is a widely used metric of system performance. Under this premise, the curve under MES consists of normal operation stages and three kinds of responses (including pre-event preventive response, during-event emergency response, and post-event restoration response). In the normal operation stage, the optimal economic operation is typically conducted from time $t_{0}$ to time $t_{1}$. After detecting the extreme event information, the preventive response will be carried out from time $t_{1}$ to time $t_{2}$, e.g., purchasing mobile electric vehicle resources. During the extreme event, the emergency response is adopted to minimize the load shedding from time $t_{2}$ to time $t_{3}$ so that the extent of system performance degradation could be alleviated (note that the system performance can be maintained for a short time due to the deployed mobile resources in the preventive response). Then, the degraded state could be observed from time $t_{3}$ to time $t_{4}$ at the end of event. From time $t_{4}$ to time $t_{5}$, the restoration response is achieved to restore critical loads [74] like hospitals and water pumps by adopting temporary measures, such as grid-island operation using remaining storage and mobile energy storage units (e.g., electric vehicles). These temporary measures restore partial or prioritized functionality, even if the underlying infrastructure (e.g., transmission lines and distribution equipment) is not yet fully repaired. After stabilizing essential services, the system proceeds to recover full infrastructure, e.g., repairing line and equipment.
@@
Similarly, after integrating hydrogen into MESs, the system performance curve should reflect both the physical and functional characteristics of hydrogen. In the preventive stage, additional mobile energy resources (e.g., HTTs for gas delivery and liquid-hydrogen tankers, as well as HFCVs/HFCBs used as mobile generators) can be deployed, so that system performance can be sustained longer in the emergency period. During the emergency response, if a grid outage coincides with a hydrogen-pipeline leak, part of critical demand can be supplied by fuel cells. Owing to pipeline linepack and pressure regulation, fuel cell stack-inlet pressure/flow conditions can be temporarily maintained, allowing the fuel cell to sustain its set power and thereby reduce load shedding as long as the available upstream pressure/flow remains above the requirements of hydrogen regulator and fuel cell stack. If the leak reduces the available supply below those requirements, hydrogen starvation occurs[116], leading to a rapid voltage drop and possible protective shutdown, which increases load shedding. Although linepack buffering may delay this adverse impact, fuel-cell power loss can worsen even after the initiating event ends once the buffer is depleted. Since more resources are available for deployment, an HMES can exhibit a higher performance level in the degraded state, and an example of this phenomenon can be seen in Fig. 7 of [117]. Utilizing mobile resources or conducting pipeline reconfiguration, the recovery trajectory of HMES can also be raised. Compared with MESs, hydrogen-related equipment failure in HMESs may require longer restoration time as shown in Fig. 5, which rely on random transportation network state and hydrogen flow dynamics as shown in Fig. 6 of [120] in the revised manuscript.
@@
Based on the above description, HMES resilience can be defined as the capacity of an HMES to withstand the disturbance from an extreme event (i.e., event progress during $t_{2}$ and $t_{3}$), and to rebuild and renew it afterwards[9]. Similarly, HMES resilience can also be described as the capacity to limit the system performance degradation and the duration of the degraded state in order to maintain critical services following an extreme event[5]. In addition to energy resilience[118], carbon resilience is also proposed in existing works, which describes an energy system’s ability to maintain or recover to its current and predicted emission level during and after shocks such as renewable intermittency, public health emergencies, or climate-related disruptions[53]. Although many definitions related to HMES resilience are proposed, they are originally borrowed from power grids and MESs, which can not truly reflect the physical and functional features of hydrogen[18][28][71, 72, 73, 74, 75]. For example, when an extreme event (e.g., ice disaster) causes simultaneous power outages and hydrogen pipeline leakage, even if no load loss is incurred due to backup sources, hydrogen fuel cells can not operate normally due to the loss of pressurized hydrogen supply. At this time, no resilience enhancement measure is required since there is no system performance degradation (i.e., load loss). In fact, hydrogen-related systems need immediate repair so that further pressure reduction can be avoided. Otherwise, once backup sources are unavailable, the whole HMES may collapse and result in larger negative social-economical impacts. Therefore, the definition of HMES resilience should incorporate any negative impacts (e.g., hydrogen availability and hydrogen-demand loss) of extreme events on hydrogen energy resources into consideration.
@@ IV-C5 CRI
Summary: In existing works, many load-based metrics have been widely adopted, which often translate into economic damage or loss. However, they often overlook the social impacts of service disruption. In practice, the number and types of customers affected are also critical indicators of resilience, especially during extreme events [90],[91]. For instance, shedding the same amount of load for a hospital and a residential building may lead to totally different consequences. The reason is that the former relies on power for life-saving operations, while the latter may experience only reduced comfort. Therefore, to develop a comprehensive resilience metric, it is necessary to combine load-based and customer-based metrics, as mentioned in [92]. In addition, existing metrics can not reflect the degraded state (e.g., HESS failure or hydrogen pipeline leaking) of HERs during an extreme event. Therefore, more comprehensive metrics are expected.
@@ V-B1 Multi-device fault model
where $\boldsymbol{u}=u_{\text{EL},e},~u_{\text{FC},e},~u_{\text{GT}}$. Moreover, $u_{\text{EL},e}$$u_{\text{FC},e}$$u_{\text{GT}}$ are binary variables related to electrolyzers, fuel cells, and a gas turbine, respectively. When there is a fault in a component, the corresponding value is zero. Otherwise, its value is one. Similar to the above multi-device fault model, NkN-k contingency model was considered in [99] for the stochastic failures of critical components (e.g., fuel cell stacks).
@@ V-B2 Temporal-spatial destructive model
Summary: Compared with stochastic single-type contingencies, stochastic multi-type contingency modeling has a stronger representation ability. Even so, the multi-device fault model still cannot capture the dynamics of contingencies related to extreme events, since the sequences and starting times of fault appearances are not considered. In fact, extreme events (e.g., hurricanes and floods) have temporal and spatial attributes, which may lead to multi-type contingencies with different sequences and starting times. In addition, some cascading contingencies incurred by cyber-physical attacks are not considered in existing works[102]. Above all, existing works rarely consider hydrogen-specific vulnerabilities. In fact, there are many physical vulnerabilities and cyber vulnerabilities in HMESs. In physical aspects, hydrogen tends to cause material embrittlement, particularly in metals such as steel, can lead to long-term degradation of pipelines, storage tanks, and connectors. These structural vulnerabilities increase the likelihood of mechanical failure under normal operations or stress conditions such as temperature fluctuations and pressure surges during extreme events. In cyber aspects, a hacker manipulating the pressure control in a hydrogen storage tank could trigger an overpressure event, potentially causing mechanical damage or explosion. Similarly, false data injection into hydrogen demand forecasts or electrolyzer schedules could destabilize coordinated operations across the electricity-hydrogen interface[60], leading to misallocation of resources or even blackouts. However, most contingency models treat hydrogen assets as idealized components with fixed performance characteristics in existing works, ignoring the risks posed by leakage, material fatigue, and cyber intrusion. Therefore, future research should develop risk-aware, vulnerability-inclusive contingency modeling that capture both the physical fragilities and cyber risks related to hydrogen.
@@ VI-A Single-type Hydrogen Facilities
Figure 7: The framework of two-stage resilience planning method
@@
Ref. Object Events and contingencies Hydrogen facility investment decisions Demand types Parameter/event uncertainty Optimization objectives Problem types, solving methods
[103] Power systems based on HESSs Deterministic impacts caused by heatwave events HESS (capacity) Electricity $\times$ Minimizing the total cost Single-level deterministic MILP, CPLEX
[28] Hydrogen-integrated airport energy system Insufficient power supply capacity caused by extreme events HESS (capacity) Electricity, hydrogen $\times$ Minimizing the overall economic cost Single-level deterministic MILP, CPLEX
[96] Integrated electricity-gas energy system with HRSs Stochastic line outages incurred by typhoons HRS (capacity, location, quantity) Electricity, gas $\times$ Minimizing the sum of investment cost and expected operation cost, minimizing operation costs under scenarios Two-stage MISOCP, Gurobi
[104] Hydrogen microgrids Deterministic line outages by natural disasters HESS (capacity, location, quantity) Electricity $\times$ Minimize the total cost; maximize the operation cost by selecting the worst extreme scenario; minimize load shedding and switch action costs Tri-level robust MILP, mixed-integer L-shaped method
[46] Electricity-hydrogen island energy system Stochastic multi-device failures and power interruptions during extreme events HESS (capacity) Electricity, hydrogen Wind power and electric load Maximizing the total profit, minimum value in the third level, maximizing profit minus load shedding cost Tri-level robust MILP, nested C&CG
[106] European electricity systems Low renewable generation during Dunkelflaute events HESS (location, capacity) Electricity Renewable generation Minimize total annualized costs; maximizes the operation cost by selecting the worst scenario; minimize the operation cost after the uncertainty realization Tri-level robust MILP; C&CG
[47] Electricity-hydrogen integrated energy system Stochastic line outages incurred by typhoons H2P (capacity, location, quantity) Electricity, hydrogen Event uncertainty Minimizing investment cost and expected worst-case system cost Tri-level stochastic robust optimization problem, nested C&CG with progressive hedging
[107] Gas-electric-hydrogen microgrids NkN-k contingencies caused by extreme events HESS (location, quantity) Electricity, heat $\times$ Minimizing the total cost, maximizing load restoration Bi-level MILP, CPLEX
@@
Figure 9: Illustration of line outages in an electricity-hydrogen integrated energy system under typhoons
@@
Ref. Object Events and contingencies Hydrogen facility investment decisions Demand types Parameter/event uncertainty Optimization objectives Problem types, solving methods
@@
[108] Integrated gas and electricity systems with HESSs Two kinds of deterministic extreme weather events HESS, hydrogen production facilities (capacity, location) Electricity, heating $\times$ Minimizing the total cost Deterministic single-level MINLP, Xpress solver
@@
[99] Hydrogen-based microgrids NkN-k stochastic device failure caused by disastrous events HESS, hydrogen-driven CCHP, SHS (capacity, location, quantity) Electricity, heating, cooling Renewable generation and demand Minimizing the total cost, minimizing the operation cost Two-stage stochastic MINLP, dual cutting-plane decomposition algorithm
@@
[109] Hydrogen-based seaport MESs NkN-k line outage contingencies Fuel cell electric truck, HRS, renewable generator, and remote control control (capacity, location) Electricity, heating, cooling Renewable generation, demand, and price Minimizing investment cost and expected operation cost Two-stage stochastic MINLP, Gurobi
@@
[110] Hydrogen-enriched power distribution networks NkN-k contingencies caused by extreme events HESS, MHER (location, capacity) Electricity, hydrogen Renewable generation, electrical and hydrogen load, and network contingencies Minimize the total cost; minimize the expected operation cost Risk-averse two-stage stochastic MILP, Gurobi
@@
[22] Hydrogen-based microgrids and hydrogen fuel cell bus NkN-k line contingencies incurred by extreme weather HRS and hydrogen fuel cell electric bus (capacity, location, quantity) Electricity, hydrogen $\times$ Maximizing profit, minimum value of maximized load restoration, maximizing load restoration Tri-level MILP, nested C&CG
[111] Electricity-hydrogen distribution networks Stochastic power line and hydrogen pipeline contingencies caused by extreme weather Hydrogen pipeline hardening, HESS (capacity) Electricity, hydrogen $\times$ Minimizing the maximum load shedding under the worst-case scenario, minimizing load shedding Distributionally robust MISOCP, C&CG algorithm
@@
[112] Hydrogen-electrical microgrids for sea islands NkN-k contingencies caused by extreme wind events HESS (location, capacity), hydrogen vessel purchasing and routing Electricity Electricity demand, renewable generation, line outages, hydrogen vessel availability Minimize total annualized costs; minimize operational and maintenance costs under the worst-case probability distribution Distributionally robust MILP, C&CG with strong cutting planes
@@
[113] Renewable-dominated power systems Stochastic line and tower damage under extreme weather Fuel cell, battery, electrolyzer, and hydrogen tank (capacity, rated power, location) Electricity, hydrogen Renewable generation and demand Minimizing configuration cost and operation cost Bi-level problem, PSO and MILP
@@
TABLE VIII: Modeling Frameworks of Planning Enhancement Problem for HMESs
@@
Single-level deterministic optimization Classical optimization with known parameters; no uncertainty or hierarchical decision structure $\min_{x\in X}c^{\top}x+f(x)$ Linear/Nonlinear/Mixed-integer programming solvers (e.g., Gurobi, CPLEX, IPOPT, Xpress)[103][108][28]
Two-stage stochastic programming First-stage “here-and-now” decisions made before uncertainty is realized; second-stage “wait-and-see” recourse decisions after uncertainty is revealed $\min_{x\in X}c^{\top}x+\mathbb{E}_{\xi}[Q(x,\xi)]$, where $Q(x,\xi)=\min_{y}\{q^{\top}y:Wy+Tx\geq h\}$ Column-and-Constraint Generation (C&CG)[111], dual cutting-plane decomposition[99], commercial solvers (e.g., Gurobi[109][110][96])
@@
Tri-level stochastic robust optimization Combines stochastic and robust elements: outer stage handles here-and-now, middle stage captures uncertainty distribution/worst-case, inner stage executes corrective recourse $\min_{x}c^{\top}x+\mathbb{E}_{\xi}[\max_{u\in\mathcal{U}(\xi)}\min_{y}f(x,y,u)]$ Nested C&CG with progressive hedging[47]
Distributionally robust optimization Decisions are robust against the worst-case probability distribution within an ambiguity set, rather than a single known distribution $\min_{x\in X}c^{\top}x+\sup_{P\in\mathcal{P}}\mathbb{E}_{\xi\sim P}[f(x,\xi)]$, where $f(x,\xi)=\min_{y\in Y(x,\xi)}\{q^{\top}y:Wy+Tx\geq h\}$ C&CG with strong cutting planes[112]

### v3

@@ IV-B Definitions of HMES Resilience
Before describing the definitions of HMES resilience, we first provide two system performance curves related to MES and HMES under extreme events in Fig. 5, respectively. Although the typical curve under MES cannot perfectly capture the realistic system performance trajectories (e.g., local performance fluctuations) as illustrated in [68, 69, 70], it can reflect the system’s ability to prepare for, absorb, and recover from disturbances and has been widely used in existing works[50],[71, 72, 73, 74, 75, 76]. In addition, while it is not tailored to hydrogen-enabled MESs, the performance trajectory can be applied to hydrogen-enabled MESs after minor modifications. By comparing the differences between two curves, the physical and functional features of hydrogen can be observed. Firstly, we introduce the whole process of resilience enhancement under MES. In existing works, load loss is a widely used metric of system performance. Under this premise, the curve under MES consists of normal operation stages and three kinds of responses (including pre-event preventive response, during-event emergency response, and post-event restoration response). In the normal operation stage, the optimal economic operation is typically conducted from time $t_{0}$ to time $t_{1}$. After detecting the extreme event information, the preventive response will be carried out from time $t_{1}$ to time $t_{2}$, e.g., purchasing mobile electric vehicle resources. During the extreme event, the emergency response is adopted to minimize the load shedding from time $t_{2}$ to time $t_{3}$ so that the extent of system performance degradation could be alleviated (note that the system performance can be maintained for a short time due to the deployed mobile resources in the preventive response). Then, the degraded state could be observed from time $t_{3}$ to time $t_{4}$ at the end of the event. From time $t_{4}$ to time $t_{5}$, the restoration response is achieved to restore critical loads [74] like hospitals and water pumps by adopting temporary measures, such as grid-island operation using remaining storage and mobile energy storage units (e.g., electric vehicles). These temporary measures restore partial or prioritized functionality, even if the underlying infrastructure (e.g., transmission lines and distribution equipment) is not yet fully repaired. After stabilizing essential services, the system proceeds to recover full infrastructure, e.g., repairing lines and equipment.
@@
Similarly, after integrating hydrogen into MESs, the system performance curve should reflect both the physical and functional characteristics of hydrogen. In the preventive stage, additional mobile energy resources (e.g., HTTs for gas delivery and liquid-hydrogen tankers, as well as HFCVs/HFCBs used as mobile generators) can be deployed, so that system performance can be sustained longer in the emergency period. During the emergency response, if a grid outage coincides with hydrogen-pipeline leakage, part of critical demand can be supplied by fuel cells. Owing to pipeline linepack and pressure regulation, fuel-cell stack inlet pressure/flow conditions can be temporarily maintained, allowing the fuel cell to sustain its set power and thereby reduce load shedding as long as the available upstream pressure/flow remains above the requirements of hydrogen regulators and fuel cell stacks. If the leakage reduces the available supply below those requirements, hydrogen starvation occurs[133], leading to a rapid voltage drop and possible protective shutdown, which increases load shedding. Although linepack buffering may delay this adverse impact, fuel-cell power output may continue to decline even after the initiating event ends once the buffer is depleted. Since more resources are available for deployment, an HMES can exhibit a higher performance level in the degraded state, and an example of this phenomenon can be seen in Fig. 7 of [134]. Utilizing mobile resources or conducting pipeline reconfiguration, the recovery trajectory of HMES can also be raised. Compared with MESs, hydrogen-related equipment failures in HMESs may involve longer restoration time, as illustrated in Fig. 5. This extended recovery is influenced by stochastic transportation network states and hydrogen flow dynamics as supported by Fig. 6 of [135].
@@
Based on the above description, HMES resilience can be defined as the capacity of an HMES to withstand the disturbance from an extreme event (i.e., event progress during $t_{2}$ and $t_{3}$), and to rebuild and renew it afterwards[9]. Similarly, HMES resilience can also be described as the capacity to limit the system performance degradation and the duration of the degraded state in order to maintain critical services following an extreme event[5]. Although many definitions related to HMES resilience are proposed, they are originally borrowed from power grids and MESs, which can not truly reflect the physical and functional features of hydrogen[18][28][71, 72, 73, 74, 75]. For example, when an extreme event (e.g., ice disaster) causes simultaneous power outages and hydrogen pipeline leakage, even if no load loss is incurred due to backup sources, hydrogen fuel cells may not operate normally due to the loss of pressurized hydrogen supply. At this time, no resilience enhancement measure is required since there is no system performance degradation (i.e., load loss). In fact, hydrogen-related systems need immediate repair so that further pressure reduction can be avoided. Otherwise, once backup sources are unavailable, the whole HMES may collapse and result in larger negative social and economic impacts. Therefore, the definition of HMES resilience should take any negative impacts (e.g., hydrogen availability and hydrogen-demand loss) of extreme events on hydrogen energy resources into consideration.
@@ IV-C5 CRI
Summary: In existing works, many load-based metrics have been widely adopted, which often translate into economic damage or loss. However, they often overlook the social impacts of service disruption. In practice, the number and types of customers affected are also critical indicators of resilience, especially during extreme events [91],[92]. For instance, shedding the same amount of load for a hospital and a residential building may lead to totally different consequences. The reason is that the former relies on power for life-saving operations, while the latter may experience only reduced comfort. Therefore, to develop a comprehensive resilience metric, it is necessary to combine load-based and customer-based metrics, as mentioned in [93]. In addition, existing metrics can not reflect the degraded state (e.g., HESS failure or hydrogen pipeline leakage) of HERs during an extreme event. Therefore, more comprehensive metrics are expected.
@@ V-B1 Multi-device fault model
where $\boldsymbol{u}=(u_{\text{EL},e},~u_{\text{FC},e},~u_{\text{GT}})$. Moreover, $u_{\text{EL},e}$$u_{\text{FC},e}$$u_{\text{GT}}$ are binary variables related to electrolyzers, fuel cells, and a gas turbine, respectively. When there is a fault in a component, the corresponding value is zero. Otherwise, its value is one. Similar to the above multi-device fault model, NkN-k contingency model was considered in [100] for the stochastic failures of critical components (e.g., fuel cell stacks).
@@ V-B2 Temporal-spatial destructive model
Summary: Compared with stochastic single-type contingencies, stochastic multi-type contingency modeling has a stronger representation ability. Even so, the multi-device fault model still cannot capture the dynamics of contingencies related to extreme events, since the sequences and starting times of fault appearances are not considered. In fact, extreme events (e.g., hurricanes and floods) have temporal and spatial attributes, which may lead to multi-type contingencies with different sequences and starting times. In addition, some cascading contingencies incurred by cyber-physical attacks are not considered in existing works[103]. Overall, existing works rarely consider hydrogen-specific vulnerabilities. In fact, there are many physical vulnerabilities and cyber vulnerabilities in HMESs. In physical aspects, hydrogen tends to cause material embrittlement, particularly in metals such as steel, which can lead to long-term degradation of pipelines, storage tanks, and connectors. These structural vulnerabilities increase the likelihood of mechanical failure under normal operations or stress conditions such as temperature fluctuations and pressure surges during extreme events. In cyber aspects, a hacker manipulating the pressure control in a hydrogen storage tank could trigger an overpressure event, potentially causing mechanical damage or explosion. Similarly, false data injection into hydrogen demand forecasts or electrolyzer schedules could destabilize coordinated operations across the electricity-hydrogen interface[60], leading to misallocation of resources or even blackouts. However, most contingency models treat hydrogen assets as idealized components with fixed performance characteristics in existing works, ignoring the risks posed by leakage, material fatigue, and cyber intrusion. Therefore, future research should develop risk-aware, vulnerability-inclusive contingency modeling that captures both the physical fragilities and cyber risks related to hydrogen.
@@ VI-A Single-type Hydrogen Facilities
Figure 7: The framework of two-stage resilience planning method[97]
@@
Ref. Parameter/event uncertainty Contingency uncertainty Uncertainty processing methods Hydrogen facility investment decisions Optimization objective components Problem types and solving methods
[104] $\times$ Deterministic impacts $\times$ HESS (capacity) Total cost Single-level deterministic MILP, CPLEX
[28] $\times$ Deterministic impacts $\times$ HESS (capacity) Overall economic cost Single-level deterministic MILP, CPLEX
[97] $\times$ Line fault uncertainty Line fault scenarios HRS (capacity, location, quantity) Investment cost and expected operation cost Two-stage MISOCP, Gurobi
[105] Natural disaster event uncertainty Component faults and fault durations Uncertainty set, natural disaster scenarios HESS (capacity, location, quantity) Total cost, operation cost, load shedding, and switching cost Tri-level robust MILP, mixed-integer L-shaped method
[46] $\times$ Fault seasons and component fault combinations Season scenarios, fault uncertainty sets HESS (capacity) Total profit, load shedding cost Tri-level robust MILP, nested C&CG
[107] $\times$ Low renewable generation Cardinality-constrained uncertainty sets HESS (location, capacity) Total annualized costs; operation cost Tri-level robust MILP; C&CG
[47] Typhoon event uncertainty Stochastic line outages incurred by typhoons Typhoon scenarios, line outage uncertainty sets H2P (capacity, location, quantity) Investment cost and expected worst-case system cost Tri-level stochastic robust optimization, nested C&CG with hedging
[108] $\times$ NkN-k contingencies Contingency scenarios HESS (location, quantity) Total cost, load restoration quantity Bi-level MILP, CPLEX
@@
Figure 9: Illustration of line outages in an electricity-hydrogen integrated energy system under typhoons[47]
@@
Ref. Parameter/event uncertainty Contingency uncertainty Uncertainty processing methods Hydrogen facility investment decisions Optimization objective components Problem types and solving methods
@@
[109] $\times$ Deterministic impacts $\times$ HESS, hydrogen production facilities (capacity, location) Total cost Single-level deterministic MINLP, Xpress solver
@@
[100] Renewable generation and demand NkN-k stochastic device failure Normal operation scenarios, emergency operational scenarios HESS, hydrogen-driven CCHP, SHS (capacity, location, quantity) Total cost, operation cost Two-stage stochastic MINLP, dual cutting-plane algorithm
@@
[110] Renewable generation, demand, and price NkN-k line outage contingencies Normal operation scenarios, emergency operational scenarios Fuel cell electric truck, HRS, renewable generator, and remote switch (capacity, location) Investment cost, expected operation cost Two-stage stochastic MINLP, Gurobi
@@
[111] Renewable generation, electrical and hydrogen load NkN-k contingencies Normal operation scenarios, network disruption scenarios HESS, MHER (location, capacity) Total cost, expected operation cost Risk-averse two-stage stochastic MILP, Gurobi
@@
[22] $\times$ NkN-k line contingencies Cardinality-constrained uncertainty sets HRS and hydrogen fuel cell electric bus (capacity, location, quantity) Profit, load restoration Tri-level MILP, nested C&CG
[112] Wind speed and rainfall intensity Power line and hydrogen pipeline uncertainties Endogenous second-order moment ambiguity set, scenarios Hydrogen pipeline hardening, HESS (capacity) Load shedding Distributionally robust MISOCP, C&CG algorithm
@@
[113] Demand, renewable generation, line outages, vessel availability NkN-k contingencies Decision dependent-based ambiguity set, scenarios HESS (location, capacity), hydrogen vessel purchasing and routing Total costs; operational and maintenance costs Distributionally robust MILP, C&CG with strong cutting planes
@@
[114] $\times$ Deterministic impacts $\times$ Fuel cell, battery, electrolyzer, and hydrogen tank (capacity, rated power, location) Configuration cost, operation cost Bi-level problem, PSO and MILP
TABLE VIII: Classification of Uncertainty Processing Methods in HMES Planning Studies
Category Key idea Typical representations Representative characteristics Formulation framework
Scenario-based methods Uncertainty is represented by a finite number of discrete scenarios, each corresponding to a possible realization of uncertain parameters, extreme events, or contingencies. Renewable generation scenarios[100]; load scenarios[110]; natural disaster scenarios[105]; line outage scenarios[47] Suitable for explicitly modeling the evolution of uncertain events; may suffer from scenario explosion and high computational burden Stochastic programming
Uncertainty-set-based methods Uncertainty is characterized by deterministic sets, and decisions are optimized to remain feasible or robust against all realizations within the prescribed sets. Interval sets; box sets; cardinality-constrained uncertainty sets[22][107] Does not require exact probability distributions; usually provides strong robustness guarantees, but may be conservative Robust optimization
Ambiguity-set-based methods Uncertainty is modeled through a family of probability distributions rather than a single known distribution, and the decision is optimized against the worst-case distribution within the ambiguity set. Moment-based ambiguity sets[112]; decision-dependent ambiguity sets[113] Suitable when the exact probability distribution is unknown but partial statistical information is available; balances robustness and probabilistic information Distributionally robust optimization
@@
TABLE IX: Modeling Frameworks and Solving Methods of Planning Enhancement Problem for HMESs
@@
Single-level deterministic optimization Classical optimization with known parameters; no uncertainty or hierarchical decision structure $\min_{x\in X}c^{\top}x+f(x)$ Linear/Nonlinear/Mixed-integer programming solvers (e.g., Gurobi, CPLEX, IPOPT, Xpress)[104]
Two-stage stochastic programming First-stage decisions made before uncertainty is realized; second-stage recourse decisions made after uncertainty is revealed $\min_{x\in X}c^{\top}x+\mathbb{E}_{\xi}[Q(x,\xi)]$, where $Q(x,\xi)=\min_{y}\{q^{\top}y:Wy+Tx\geq h\}$ Column-and-Constraint Generation (C&CG)[112], dual cutting-plane decomposition[100], commercial solvers (e.g., Gurobi[110][111])
@@
Tri-level stochastic robust optimization Outer stage handles here-and-now, middle stage captures uncertainty distribution/worst-case, inner stage executes corrective recourse $\min_{x}c^{\top}x+\mathbb{E}_{\xi}[\max_{u\in\mathcal{U}(\xi)}\min_{y}f(x,y,u)]$ Nested C&CG with progressive hedging[47]
Distributionally robust optimization Decisions are robust against the worst-case probability distribution within an ambiguity set, rather than a single known distribution $\min_{x\in X}c^{\top}x+\sup_{P\in\mathcal{P}}\mathbb{E}_{\xi\sim P}[f(x,\xi)]$ s.t. $f(x,\xi)=\min_{y\in Y(x,\xi)}\{q^{\top}y:Wy+Tx\geq h\}$ C&CG with strong cutting planes[113]